On an xy-coordinate plane, line "L" contains the points (0,0) and (1,2). If line "M" contains the point (0,0) and is perpendicular to line "L", what is an equation of line "M"? Thanks so much!
Answer by venugopalramana(3286)

On an xy-coordinate plane, line "L" contains the points (0,0) and (1,2). If line "M" contains the point (0,0) and is perpendicular to line "L", what is an equation of line "M"? Thanks so much!
SLOPE OF LINE L IS =(2-0)/(1-0)=2
SLOPE OF LINE PERPENDICULAR TO IT IS -1/2
EQN .OF LINE IS
Y-0=(-1/2)(X-0)
2Y+X=0